# A Polarization Hall Effect in Hydrated DNA

### v1

@@ III Discussion
Earlier reports of staircase-like structures and oscillations [Pietruszka2025a, Pietruszka2025b] remain quantitatively valid; the present work reframes their interpretation while preserving the underlying observations. This perspective is consistent with broader suggestions that biological systems may operate in regimes where coherence or quantum-like vibrational organization plays a functional role [AlKhaliliMcFadden2015]. Within this framework, the emergence of quantized plateaus and SdH-like oscillations follows naturally from the quasi-two-dimensional confinement of polarization modes within the hydrated DNA layer.
@@
Recasting the observed response in terms of polarization currents resolves conceptual difficulties inherent to charge-transport models. No measurable steady current flows through the sample, yet exhibits a robust structure that scales reproducibly with the magnetic field. This behavior is consistent with a macroscopic polarization mechanism, in which oscillating dipoles embedded in a constrained hydrogen-bond network generate time-dependent electric fields detected as . Within this framework, the observed plateaus correspond to stable configurations of coherent dipolar order, and their spacing reflects the quantization of dipolar orientation states under magnetic torque.
This interpretation is consistent with the microscopic mechanism illustrated in Fig. 1. When a proton transiently hops along a hydrogen bond, it leaves behind a correlated vacancy, forming a locally neutral proton–proton-hole pair. Although electrically neutral, such dipoles are strongly polarizable and can acquire an effective magnetic torque arising from the combination of orbital motion and hydrogen-bond geometry. When many dipoles synchronize through the hydration network, the resulting collective dynamics produce a macroscopic, time-dependent polarization , whose time derivative gives rise to the observed transverse voltage . This collective polarization mode fits naturally within the framework of Fröhlich-like coherence, corresponding to long-range phase ordering in a driven, dissipative system [Frohlich1968, Frohlich1988, Davydov1979].
In this picture, the magnetic field couples to the dipolar ensemble through an effective magnetoelectric interaction. Field-induced torques bias the orientation and phase of proton-proton-hole dipoles, while their collective reorientation feeds back into the macroscopic polarization response. A detailed microscopic magnetoelectric tensor is not required at this stage; it is sufficient that the chiral hydrogen-bond geometry and finite loop areas allow magnetic flux to modulate the phase of the coherent dipolar mode. This mechanism provides a natural route by which moderate magnetic fields can reshape the polarization landscape and enable discrete, quantized reconfigurations of the dipolar ensemble.
@@
The staircase structures observed in and the periodic oscillations in represent two complementary manifestations of the same underlying quantization process. At lower magnetic fields, the dipolar ensemble subdivides into domains that realign sequentially as the field increases, producing SdH-like oscillations in both and . The associated frequencies reflect characteristic coherence areas defined by the structured hydration layer. As the field approaches the critical value near T, these domains merge and the system undergoes a sharp transition into a Landau-like regime characterized by quantized polarization plateaus, which constitute the macroscopic signature of field-locked dipolar coherence.
In this regime, the polarization modes can be viewed as quantized angular-momentum states of a collective dipolar field. The quantization arises from coupling between magnetic flux and the effective areas of coherently oscillating dipolar loops, analogous to Onsager-type relations in electronic systems. The extracted coherence densities, consistently in the range – across SdH-like oscillations, plateau spacing, and tensor inversion, support the presence of a single underlying coherence ensemble. These values correspond to plausible surface coverages of hydrogen-bonded interfacial water, implying coherence domains on the order of several hundred nanometers, in agreement with recent structural studies of confined hydration layers [DonadioGalli2025].
@@
Above the threshold magnetic field, the oscillatory structure collapses into broad, slowly varying lobes, indicating a qualitative reorganization of the dipolar ensemble. The discrete domain architecture gives way to a globally phase-locked state in which a large fraction of dipoles oscillate coherently. This Fröhlich-like phase is consistent with a regime in which magnetic alignment suppresses dominant decoherence pathways, allowing energy injected into local modes to accumulate into a single, collective oscillation [Pietruszka2025b]. The concomitant flattening of the transverse response () and the smooth saturation of the longitudinal component () above the threshold are consistent with a crossover from a domain-based regime to a globally coherent polarization mode.
Such behavior mirrors transitions discussed in classical Fröhlich systems and bears qualitative resemblance to aspects of Bose–Einstein condensation [Anderson1995, Ketterle1999], although realized here in a driven, dissipative biological matrix at ambient conditions. The persistence of coherence even when substantially exceeds the estimated dipolar coupling energies suggests a nontrivial, self-organized mechanism facilitated by nonlinear protonic interactions within the hydrogen-bond network.
@@
The temperature dependence of the polarization oscillations reveals a coherence boundary near K, close to biologically relevant temperatures [PietruszkaMarzec2024]. With increasing temperature, thermal motion progressively disrupts dipolar phase locking, leading to a suppression of quantized features. Cooling reverses this trend, first restoring discrete oscillations and ultimately yielding fully quantized polarization states at lower temperatures. The amplitudes of the oscillatory components extracted from FFT analysis decay approximately as , with activation energies comparable to hydrogen-bond strengths, directly linking the emergence of coherence to the structured hydration network. The appearance of a coherence maximum near physiological temperature suggests that biological hydration layers may naturally operate in a regime optimized by the balance between thermal agitation and collective ordering.
@@
Additional evidence that the magnetic-field-driven transition reflects a collective reorganization of the hydrated DNA system is provided by the photovoltaic response shown in Fig. 12. A qualitatively similar transition appears in this independent proxy channel, which is spatially separated from the electrical contacts. The correspondence between electrical and photovoltaic signatures indicates that the field-induced transition involves a redistribution of polarization dynamics across the DNA–water interface, capable of coupling to both electrical and electromagnetic observables. Although the microscopic origin of the photovoltaic response remains to be clarified, its synchronization with the electrical transition supports the interpretation of a system-level coherent polarization mode.
@@
The findings reported here provide a unified interpretation of several years of experimental observations on hydrated DNA under magnetic fields [Pietruszka2025a, Pietruszka2025b]. Although earlier analyses framed the transverse response in terms of protonic conduction, all quantitative features—critical fields, SdH-like periodicities, plateau spacings, threshold voltages, and inferred coherence densities—remain internally consistent across studies. The present work reinterprets these features as signatures of field-organized dipolar polarization rather than transport by mobile charge carriers. The recurrence of the effective coherence density across independent datasets, including those reported in Physica B and BioSystems, underscores the robustness of the underlying physical phenomenon.
@@
The interference pattern shown in Fig. 10 provides direct evidence of coherent coupling between the longitudinal and transverse polarization channels. By combining the simultaneously acquired signals and time-shifted into the interference map , we observe alternating regions of enhanced and suppressed amplitude that depend sensitively on their relative phase. These structured bands arise from the superposition of two phase-locked oscillatory components and reflect the instantaneous phase relationship between the longitudinal and transverse polarization dynamics.
@@
Transverse responses carried by neutral excitations have become a central theme in modern condensed-matter physics. Recent demonstrations of the phonon Hall effect, including a recent Phys. Rev. Lett. report [Jin2025], show that neutral vibrational excitations can generate measurable transverse voltages in insulating crystals. This growing body of work, encompassing phonon thermal Hall effects [Strohm2005, Grissonnanche2019, ZhangEtAl2023, Xiang2025], magnon Hall currents [Onose2010], and exciton Hall responses [Mak2018], establishes that Berry curvature, chiral geometry, and broken inversion symmetry can produce Hall-like phenomena in the absence of charge transport.
@@
Viewed through this lens, the polarization Hall effect reflects a broader principle: soft, hydrogen-bonded matter can support magnetically induced transverse responses mediated by coherent polarization waves. This perspective points to an emerging area of study—protonic polarization physics—focused on the coherence, geometry, and controllability of proton-mediated polarization modes in aqueous and biological environments. Related concepts may connect the dynamics of hydrated DNA to coherent energy-transfer phenomena observed in photosynthetic complexes [Engel2007, Panitchayangkoon2010] and to other forms of water-mediated collective behavior.
@@
IV Conclusion
The experiments reported here establish hydrated DNA as a field-stabilized and temperature-gated coherent dipolar system. Moderate magnetic fields and controlled cooling reorganize the hydration layer into discrete, quantized polarization states and, at lower temperatures, into robust oscillatory modes. The observed staircase-like transitions, well-resolved -periodic oscillations, and large-amplitude transverse responses are not consistent with conventional ionic or electronic transport. Instead, they arise from the collective dynamics of proton–proton-hole dipoles within the hydrogen-bond network, whose synchronized motion generates a measurable polarization current.
@@ Methods
Barley genomic DNA (double-stranded, isolated using the CTAB method, modified after Doyle and Doyle [DoyleDoyle1987]) was dissolved in TE buffer (pH 8, containing Tris–HCl and EDTA). Working concentrations were 100, 500, and 1000 ng L-1. Prepared DNA solutions were stored at C between experiments to minimize degradation.
@@
Magnetic fields in the range 0.05–0.65 T were generated using NdFeB permanent magnets mounted on a motor-driven linear translation stage that controlled the magnet–sample separation. The system operated in two modes: a motor-sweep mode (0.1–0.5 mT s-1) for high-resolution oscillation and plateau measurements, and a manual mode for threshold measurements to minimize potential electromagnetic interference from the motor. Magnetic-field values were calibrated before each measurement session using a Hall-probe gaussmeter with an accuracy of mT. All fields were applied perpendicular to the sample plane.
@@
Magnetic-field sweep experiment shown in Fig. 4 was sampled at 740 Hz, yielding approximately data points per trace. The remaining magnetic-field and temperature sweeps were sampled at 1 Hz. All measurement electronics were battery-powered and magnetically shielded to minimize inductive coupling.
@@
Extraction of effective coherence area and density
The periodicity in was converted into a characteristic dipolar coherence area using an Onsager-type geometric estimator,
The appearance of the elementary charge in this Onsager-type estimator reflects its standard geometric form and does not imply electronic carriers in the present polarization-based interpretation.
This expression was used solely to estimate the characteristic spatial scale of coherent dipolar domains and not to infer electronic carrier density. Independent estimates derived from FFT periodicity, plateau spacing, and comparisons across measurement protocols consistently yielded –, indicating a single field-organized dipolar coherence ensemble.
@@
The magnetic-field–temperature (–) phase diagram was constructed using magnetic-field thresholds near 0.25–0.35 T, temperature transitions at 20.6 ∘C and 12.0 ∘C, polarization plateau structures, SdH-like oscillation bands, and the onset of low-temperature Fröhlich-type oscillations. Gaussian blurring was applied to avoid artificially sharp boundaries, yielding a continuous representation of coherent dipolar regimes.
@@
All data supporting the findings of this study are openly available. The complete transverse voltage dataset for Fig. 4 (approximately 100,000 data points) is deposited on Zenodo and can be accessed via the DOI 10.5281/zenodo/1212_121326_8. Additional raw data, processed data, and plotting scripts are available from the corresponding author upon reasonable request.

### v2

@@ III Discussion
Earlier reports of staircase-like structures and oscillations [21, 20] remain quantitatively valid; the present work reframes their interpretation while preserving the underlying observations. This perspective is consistent with broader suggestions that biological systems may operate in regimes where coherence or quantum-like vibrational organization plays a functional role [1]. Within this framework, the emergence of quantized plateaus and -periodic oscillations follows naturally from the quasi-two-dimensional confinement of polarization modes within the hydrated DNA layer.
@@
Recasting the observed response in terms of polarization currents resolves conceptual difficulties inherent to charge-transport models. No measurable steady current flows through the sample, yet exhibits a robust structure that scales reproducibly with the magnetic field. This behavior is consistent with a macroscopic polarization mechanism, in which oscillating dipoles embedded in a constrained hydrogen-bond network generate time-dependent electric fields detected as . Within this framework, the observed plateaus correspond to stable configurations of coherent dipolar order, and their spacing reflects the quantization of dipolar orientation states under magnetic torque. From a dynamical perspective, these plateaus correspond to metastable attractors of a driven polarization phase field. Transitions between neighboring plateaus reflect switching events induced by magnetic-field tuning and assisted by thermal and environmental fluctuations.
This interpretation is consistent with the microscopic mechanism illustrated in Figure 1. When a proton transiently hops along a hydrogen bond, it leaves behind a correlated vacancy, forming a locally neutral proton–proton-hole pair. Although electrically neutral, such dipoles are strongly polarizable and can acquire an effective magnetic torque arising from the combination of orbital motion and hydrogen-bond geometry. When many dipoles synchronize through the hydration network, the resulting collective dynamics produce a macroscopic, time-dependent polarization , whose time derivative gives rise to the observed transverse voltage . This collective polarization mode fits naturally within the framework of Fröhlich-like coherence, corresponding to long-range phase ordering in a driven, dissipative system [7, 8, 3].
In this picture, the magnetic field couples to the dipolar ensemble through an effective magnetoelectric interaction. Field-induced torques bias the orientation and phase of proton-proton-hole dipoles, while their collective reorientation feeds back into the macroscopic polarization response. A detailed microscopic magnetoelectric tensor is not required at this stage; it is sufficient that the chiral hydrogen-bond geometry and finite loop areas allow magnetic flux to influence the phase of the coherent dipolar mode. This mechanism provides a natural route by which moderate magnetic fields can reshape the polarization landscape and enable discrete, quantized reconfigurations of the dipolar ensemble.
@@
The staircase structures observed in and the periodic oscillations in represent two complementary manifestations of the same underlying quantization process. At lower magnetic fields, the dipolar ensemble subdivides into domains that realign sequentially as the field increases, producing -periodic oscillations in both and . The associated frequencies reflect characteristic coherence areas defined by the structured hydration layer. As the field approaches the critical value near T, these domains merge and the system undergoes a sharp transition into a Landau-like regime characterized by quantized polarization plateaus, which constitute the macroscopic signature of field-locked dipolar coherence.
In this regime, the polarization modes can be viewed as quantized angular-momentum states of a collective dipolar field. The quantization arises from coupling between magnetic flux and the effective areas of coherently oscillating dipolar loops, analogous to Onsager-type relations in electronic systems. The extracted coherence densities, consistently in the range – across -periodic oscillations, plateau spacing, and tensor inversion, support the presence of a single underlying coherence ensemble. These values correspond to plausible surface coverages of hydrogen-bonded interfacial water, implying coherence domains on the order of several hundred nanometers, in agreement with recent structural studies of confined hydration layers [4].
@@
Above the threshold magnetic field, the oscillatory structure collapses into broad, slowly varying lobes, indicating a qualitative reorganization of the dipolar ensemble. The discrete domain architecture gives way to a globally phase-locked state in which a large fraction of dipoles oscillate coherently. This Fröhlich-like phase is consistent with a regime in which magnetic alignment suppresses dominant decoherence pathways, allowing energy injected into local modes to accumulate into a single, collective oscillation [20]. The concomitant flattening of the transverse response () and the smooth saturation of the longitudinal component () above the threshold are consistent with a crossover from a domain-based regime to a globally coherent polarization mode.
Such behavior mirrors transitions discussed in classical Fröhlich systems and bears qualitative resemblance to aspects of Bose–Einstein condensation [2, 12], although realized here in a driven, dissipative biological matrix at ambient conditions. The persistence of coherence even when substantially exceeds the estimated dipolar coupling energies suggests a nontrivial, self-organized mechanism facilitated by nonlinear protonic interactions within the hydrogen-bond network.
@@
The temperature dependence of the polarization oscillations reveals a coherence boundary near K, close to biologically relevant temperatures [19]. With increasing temperature, thermal motion progressively disrupts dipolar phase locking, leading to a suppression of quantized features. Cooling reverses this trend, first restoring discrete oscillations and ultimately yielding fully quantized polarization states at lower temperatures. The amplitudes of the oscillatory components extracted from FFT analysis decay approximately as , with activation energies comparable to hydrogen-bond strengths, directly linking the emergence of coherence to the structured hydration network. The appearance of a coherence maximum near physiological temperature suggests that biological hydration layers may naturally operate in a regime optimized by the balance between thermal agitation and collective ordering.
@@
Additional evidence that the magnetic-field-driven transition reflects a collective reorganization of the hydrated DNA system is provided by the photovoltaic response shown in Figure 12. A qualitatively similar transition appears in this independent proxy channel, which is spatially separated from the electrical contacts. The correspondence between electrical and photovoltaic signatures indicates that the field-induced transition involves a redistribution of polarization dynamics across the DNA–water interface, capable of coupling to both electrical and electromagnetic observables. Although the microscopic origin of the photovoltaic response remains to be clarified, its synchronization with the electrical transition supports the interpretation of a system-level coherent polarization mode.
@@
The findings reported here provide a unified interpretation of several years of experimental observations on hydrated DNA under magnetic fields [21, 20]. Although earlier analyses framed the transverse response in terms of protonic conduction, all quantitative features—critical fields, -periodic periodicities, plateau spacings, threshold voltages, and inferred coherence densities—remain internally consistent across studies. The present work reinterprets these features as signatures of field-organized dipolar polarization rather than transport by mobile charge carriers. The recurrence of the effective coherence density across independent datasets, including those reported in Physica B and BioSystems, underscores the robustness of the underlying physical phenomenon.
@@
The interference pattern shown in Figure 10 provides direct evidence of coherent coupling between the longitudinal and transverse polarization channels. By combining the simultaneously acquired signals and time-shifted into the interference map , we observe alternating regions of enhanced and suppressed amplitude that depend sensitively on their relative phase. These structured bands arise from the superposition of two phase-locked oscillatory components and reflect the instantaneous phase relationship between the longitudinal and transverse polarization dynamics.
@@
Transverse responses carried by neutral excitations have become a central theme in modern condensed-matter physics. Recent demonstrations of the phonon Hall effect, including a recent Phys. Rev. Lett. report [11], show that neutral vibrational excitations can generate measurable transverse voltages in insulating crystals. This growing body of work, encompassing phonon thermal Hall effects [25, 9, 31, 30], magnon Hall currents [15], and exciton Hall responses [14], establishes that Berry curvature, chiral geometry, and broken inversion symmetry can produce Hall-like phenomena in the absence of charge transport.
@@
Viewed through this lens, the polarization Hall effect reflects a broader principle: soft, hydrogen-bonded matter can support magnetically induced transverse responses mediated by coherent polarization waves. This perspective points to an emerging area of study—protonic polarization physics—focused on the coherence, geometry, and controllability of proton-mediated polarization modes in aqueous and biological environments. Related concepts may connect the dynamics of hydrated DNA to coherent energy-transfer phenomena observed in photosynthetic complexes [6, 17] and to other forms of water-mediated collective behavior.
@@
IV Minimal theoretical description
The observed transverse polarization response can be described at a minimal level as a driven, dissipative collective phase mode rather than as a transport Hall effect. We denote by a coarse-grained phase (or orientation angle) characterizing the collective polarization state of the quasi-two-dimensional hydrated DNA layer. The experimentally measured transverse voltage is treated phenomenologically as being proportional to the time derivative of the transverse polarization component,
The dynamics of are assumed to follow an overdamped evolution in an effective field- and temperature-dependent landscape,
where is a damping coefficient and represents thermal and environmental fluctuations. Discrete polarization plateaus correspond to intervals in which the system remains trapped near stable minima of , while step transitions arise when changes in magnetic field or temperature destabilize a given minimum and the system switches to a neighboring one.
Within this framework, enhanced fluctuations localized at plateau boundaries reflect competition between neighboring polarization configurations near stability thresholds rather than measurement noise. The approximately regular spacing of plateau centers and the appearance of -periodic structure preceding plateau formation are captured phenomenologically by the magnetic-field dependence of , without invoking itinerant charge carriers or electronic Landau quantization. This minimal description is intended to capture the essential features of multistability, switching, and coherence observed experimentally, while remaining agnostic about microscopic details.
V Conclusion
The experiments reported here establish hydrated DNA as a field-stabilized and temperature-gated coherent dipolar system. Moderate magnetic fields and controlled cooling reorganize the hydration layer into discrete, reproducible polarization states and, at lower temperatures, into robust oscillatory modes. The observed staircase-like transitions, well-resolved -periodic oscillations, and large-amplitude transverse responses are not consistent with conventional ionic or electronic transport. Instead, they arise from the collective dynamics of proton–proton-hole dipoles within the hydrogen-bond network, whose synchronized motion generates a measurable polarization current.
@@ Methods
Barley genomic DNA (double-stranded, isolated using the CTAB method, modified after Doyle and Doyle [5]) was dissolved in TE buffer (pH 8, containing Tris–HCl and EDTA). Working concentrations were 100, 500, and 1000 ng L-1. Prepared DNA solutions were stored at C between experiments to minimize degradation.
@@
Magnetic fields in the range 0.05–0.65 T were generated using NdFeB permanent magnets mounted on a motor-driven linear translation stage that controlled the magnet–sample separation (Figure 14). The system operated in two modes: a motor-sweep mode (0.1–0.5 mT s-1) for high-resolution oscillation and plateau measurements, and a manual mode for threshold measurements to minimize potential electromagnetic interference from the motor. Magnetic-field values were calibrated before each measurement session using a Hall-probe gaussmeter with an accuracy of mT. All fields were applied perpendicular to the sample plane.
@@
The magnetic-field sweep experiment shown in Figure 4 was sampled at 740 Hz, yielding approximately data points per trace. The remaining magnetic-field and temperature sweeps were sampled at 1 Hz. All measurement electronics were battery-powered and magnetically shielded to minimize inductive coupling.
@@
Extraction of characteristic periodicity scale
The oscillatory structures observed in the transverse and longitudinal polarization channels are periodic in . We characterize this behavior by extracting a dominant frequency scale from fast Fourier transform (FFT) analysis performed in the domain. This frequency provides a compact and reproducible descriptor of the magnetic-field-dependent reorganization of the polarization response.
For comparison across datasets and experimental protocols, we associate with an inverse-area scale defined phenomenologically as
where is a fixed proportionality constant used solely to express the periodicity in area-like units. In the present polarization-based interpretation, does not represent a Fermi-surface area or any electronic carrier property. Instead, it serves as a convenient measure of a characteristic coherence-patterning scale within the quasi-two-dimensional hydrated layer.
An associated effective density scale is defined from for the purpose of comparing oscillatory periodicities with the spacing of discrete polarization plateaus. Both quantities are treated as phenomenological descriptors of the field-organized collective polarization dynamics. The numerical values quoted in the main text are therefore used only to demonstrate internal consistency across independent measurements, and not to imply itinerant charge transport or electronic quantization.
@@
The magnetic-field–temperature (–) phase diagram was constructed using magnetic-field thresholds near 0.25–0.35 T, temperature transitions at 20.6 ∘C and 12.0 ∘C, polarization plateau structures, -periodic oscillation bands, and the onset of low-temperature Fröhlich-type oscillations. Gaussian blurring was applied to avoid artificially sharp boundaries, yielding a continuous representation of coherent dipolar regimes.
@@
All data supporting the findings of this study are openly available. The complete transverse voltage dataset for Figure 4 (approximately 100,000 data points) is deposited on Zenodo and can be accessed via the DOI 10.5281/zenodo/1212_121326_8. Additional raw data, processed data, and plotting scripts are available from the corresponding author upon reasonable request.